# Risk Bounds for Mixture Density Estimation on Compact Domains via the $h$-Lifted Kullback–Leibler Divergence

### v1

@@ Abstract
We consider the problem of estimating probability density functions based on sample data, using a finite mixture of densities from some component class. To this end, we introduce the $h$-lifted Kullback–Leibler (KL) divergence as a generalization of the standard KL divergence and a criterion for conducting risk minimization. Under a compact support assumption, we prove an $\mathcal{O}(1/{\sqrt{n}})$ bound on the expected estimation error when using the $h$-lifted KL divergence, which extends the results of Rakhlin et al. (2005, ESAIM: Probability and Statistics, Vol. 9) and Li and Barron (1999, Advances in Neural Information Processing Systems, Vol. 12) to permit the risk bounding of density functions that are not strictly positive. We develop a procedure for the computation of the corresponding maximum $h$-lifted likelihood estimators ($h$-MLLEs) using the Majorization-Maximization framework and provide experimental results in support of our theoretical bounds.
Keywords: Kullback–Leibler divergence, mixture density estimation, maximum likelihood, convergence rates, learning theory.
@@ 1 Introduction
Let $\left(\Omega,\mathfrak{A},{\mathrm{\mathbf{P}}}\right)$ be an abstract probability space and let $X:\Omega\to\mathcal{X}$ be a random variable taking values in the measurable space $\left(\mathcal{X},\mathfrak{F}\right)$, where $\mathcal{X}$ is a compact metric space equipped with its Borel $\sigma$-algebra $\mathfrak{F}$. Suppose that we observe an independent and identically distributed (i.i.d.) sample of random variables $\mathbf{X}_{n}=\left(X_{i}\right)_{i\in\left[n\right]}$, where $\left[n\right]=\left\{1,\ldots,n\right\}$, and that each $X_{i}$ arises from the same data generating process as $X$, characterized by the probability measure $F\ll\mu$ on $\left(\mathcal{X},\mathfrak{F}\right)$, with density function $f=\mathrm{d}F/\mathrm{d}\mu$, for some $\sigma$-finite $\mu$.
In this work, we are concerned with the estimating $f$ via a data dependent double-index sequence of estimators $\left(f_{k,n}\right)_{k,n\in\mathbb{N}}$, where
$f_{k,n}\in\mathcal{C}_{k}$ $=\mathrm{co}_{k}\left(\mathcal{P}\right)$
=$=\left\{f_{k}\left(\cdot;\psi_{k}\right)=\sum_{j=1}^{k}\pi_{j}\varphi\left(\cdot;\theta_{j}\right)\mid\varphi\left(\cdot;\theta_{j}\right)\in\mathcal{P},\,\pi_{j}\geq 0,\,j\in\left[k\right],\,\sum_{j=1}^{k}\pi_{j}=1\right\},$
@@
$\mathcal{P}=\left\{\varphi\left(\cdot;\theta\right):\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}\mid\theta\in\Theta\subset\mathbb{R}^{d}\right\},$
$\psi_{k}=\left(\pi_{1},\dots,\pi_{k},\theta_{1},\dots,\theta_{k}\right)$, and $d\in\mathbb{N}$. To ensure the measurability and existence of various optima, we shall assume that $\varphi$ is Caratheodory in the sense that $\varphi\left(\cdot;\theta\right)$ is $\left(\mathcal{X},\mathfrak{F}\right)$-measurable, for each $\theta\in\Theta$, and $\varphi\left(X;\cdot\right)$ is continuous for each $X\in\mathcal{X}$.
In the definition above, we can identify the set $\mathcal{C}_{k}=\mathrm{co}_{k}\left(\mathcal{P}\right)$ as the set of density functions that can be written as a convex combination of $k$ elements of $\mathcal{P}$, where $\mathcal{P}$ is often called the space of component density functions. We then interpret $\mathcal{C}_{k}$ as the class of $k$-component finite mixtures of densities of class $\mathcal{P}$, as studied, for example, by McLachlan and Peel (2004); Nguyen et al. (2020, 2022b).
@@ 1.1 Risk bounds for mixture density estimation
where $\left(p,q\right)\mapsto\ell\left(p,q\right)\in\mathbb{R}_{\geq 0}$ is a loss function on pairs of density functions. We also define the density-to-class loss
@@
where $\mathrm{cl}(\cdot)$ is the closure. Here, we identify $\left(k,n\right)\mapsto\rho\left(k,n\right)$ as a characterization of the rate at which the left-hand side of (1) converges to zero as $k$ and $n$ increase. Our present work follows the research of Li and Barron (1999), Rakhlin et al. (2005) and Klemelä (2007) (see also Klemelä 2009, Ch. 19). In Li and Barron (1999) and Rakhlin et al. (2005), the authors consider the case where $\ell\left(p,q\right)$ is taken to be the Kullback–Leibler (KL) divergence
$\mathrm{KL}\left(p\,||\,q\right)=\int p\log\frac{p}{q}\mathrm{d}\mu,$
@@
Under the assumption that $f,f_{k}\geq a$, for some $a>0$ and each $k\in\left[n\right]$ (i.e., strict positivity), Li and Barron (1999) obtained the bound
@@
for constants $c_{1},c_{2},c_{3}>0$, which was then improved by Rakhlin et al. (2005) who obtain the bound
@@
where $\lVert p\rVert_{2,\mu}^{2}=\int_{\mathcal{X}}\lvert p\rvert^{2}\mathrm{d}\mu$, for each $p\in L_{2}\left(\mu\right)$, and choose $f_{k,n}$ as minimizers of the $L_{2}\left(\mu\right)$ empirical risk: i.e., $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$, where
@@
$c_{1},c_{2}>0$, without the lower bound assumption on $f,f_{k}$ above, even permitting $\mathcal{X}$ to be unbounded. Via the main results of Naito and Eguchi (2013), the bound above can be generalized to the $U$-divergences, which includes the special $L_{2}(\mu)$ norm distance as a special case.
On the one hand, the sequence of MLEs required for the results of Li and Barron (1999) and Rakhlin et al. (2005) are typically computable, for example, via the usual expectation–maximization approach (cf. McLachlan and Peel 2004, Ch. 2). This contrasts with the computation of least-squares density estimators of form (2), which requires evaluations of the typically intractable integral expressions: $\left\|f_{k}\left(\cdot;\psi_{k}\right)\right\|_{2}^{2}$. However, the least-squares approach of Klemelä (2007) permits the analysis using families $\mathcal{P}$ of usual interest, such as normal distributions and beta distributions, the latter of which being compactly supported but having densities that cannot be bounded away from zero without restrictions, and thus do not satisfy the regularity conditions of Li and Barron (1999) and Rakhlin et al. (2005).
@@ 1.2 Main contributions
We propose the following $h$-lifted KL divergence, as a generalization of the standard KL divergence to address the computationally tractable estimation of density functions which do not satisfy the regularity conditions of Li and Barron (1999) and Rakhlin et al. (2005). The use of the $h$-lifted KL divergence has the possibility to advance theories based on the standard KL divergence in statistical machine learning. To this end, let $h:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ be a function in $L_{1}(\mu)$, and define the $h$-lifted KL divergence by:
@@
Assume that $h$ is a probability density function, and let $\mathbf{Y}_{n}=\left(Y_{i}\right)_{i\in\left[n\right]}$ be a an i.i.d. sample, independent of $\mathbf{X}_{n}$, where each $Y_{i}:\Omega\rightarrow\mathcal{X}$ is a random variable with probability measure on $(\mathcal{X},\mathfrak{F})$, characterized by the density $h$ with respect to $\mu$. Then, for each $k$ and $n$, let $f_{k,n}$ be defined via the maximum $h$-lifted likelihood estimator ($h$-MLLE): $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$, where
@@
This result is a compromise between the works of Li and Barron (1999) and Rakhlin et al. (2005), and Klemelä (2007), as it applies to a broader space of component densities $\mathcal{P}$, and because the required $h$-MLLEs, (4), can be efficiently computed via minorization–maximization (MM) algorithms (see, e.g., Lange 2016). We shall discuss this assertion in Section 6.
@@ 1.3 Relevant literature
Our work largely follows the approach of Li and Barron (1999), which have was extended upon by Rakhlin et al. (2005) and Klemelä (2007). All three texts use approaches based on the availability of greedy algorithms for maximizing convex functions with convex functional domains. In this work, we shall make use of the proof techniques of Zhang (2003). Related results in this direction can be found in DeVore and Temlyakov (2016) and Temlyakov (2016). Making the same boundedness assumption as Rakhlin et al. (2005), Dalalyan and Sebbar (2018) obtain refined oracle inequalities under the additional assumption that the class $\mathcal{P}$ is finite. Numerical implementations of greedy algorithms for estimating finite mixtures of Gaussian densities were studied by Vlassis and Likas (2002) and Verbeek et al. (2003).
The $h$-MLLE as an optimization objective can be compared to other similar modified likelihood estimators, such as the $L_{q}$ likelihood of Ferrari and Yang (2010) and Qin and Priebe (2013), the $\beta$-likelihood of Basu et al. (1998) and Fujisawa and Eguchi (2006), penalized likelihood estimators, such as maximum a posteriori estimators of Bayesian models, or $f$-separable Bregman distortion measures of Kobayashi and Watanabe (2024, 2021).
The practical computation of the $h$-MLLEs, (4), is made possible via the MM algorithm framework of Lange (2016), see also Hunter and Lange (2004), Wu and Lange (2010), and Nguyen (2017) for further details. Such algorithms have well-studied global convergence properties and can be modified for mini-batch and stochastic settings (see, e.g., Razaviyayn et al., 2013 and Nguyen et al., 2022a).
A related and popular setting of investigations is that of model selection, where the objects of interest are single-index sequences $\left(f_{k_{n},n}\right)_{n\in\mathbb{N}}$, and where the aim is to obtain finite-sample bounds for losses of the form $\ell\left(f_{k_{n},n},f\right)$, where each $k_{n}\in\mathbb{N}$ is a data dependent function, often obtained by optimizing some penalized loss criterion, as described in Massart (2007), Koltchinskii (2011, Ch. 6), and Giraud (2021, Ch. 2). In the context of finite mixtures, examples of such analyses can be found in the works of Maugis and Michel (2011) and Maugis-Rabusseau and Michel (2013). A comprehensive bibliography of model selection results for finite mixtures and related statistical models can be found in Nguyen et al. (2022c).
@@ 1.4 Organization of paper
The manuscript proceeds as follows. In the following section we formally characterize the $h$-lifted KL divergence as a Bregman divergence, and establish some of its properties. In Section 4, we prove new risk bounds of the form (1) in terms of the $h$-lifted KL divergence. We discuss the computation of the $h$-lifted likelihood estimator of form (4) in Section 6, followed by the presentation of some empirical results regarding the convergence of (1) in terms of both $k$ and $n$. Additional technical results are also included in the Appendices.
@@ 2.1 $\textrm{KL}_{h}$ as a Bregman divergence
The properties for Bregman divergences between scalars can be extended to density functions and other functional spaces, as established in Frigyik et al. (2008) and Stummer and Vajda (2012), for example. We also direct the interested reader to the works of Pardo (2006), Basu et al. (2011), and Amari (2016).
@@
The $h$-lifted KL divergence $\mathrm{KL}_{h}$ as a Bregman divergence is generated by the function $\phi(u)=(u+h)\log(u+h)-(u+h)+1$. This assertion is demonstrated in Appendix A.
@@ 2.2 Advantages of the $h$-lifted KL divergence
When the standard KL divergence is employed in the density estimation problem, it is common to restrict consideration of density functions to those bounded away from zero by some positive constant. That is, one typically considers the smaller class of so-called admissible target densities $\mathcal{P}_{\alpha}\subset\mathcal{P}$ (cf. Meir and Zeevi, 1997), where
@@
$\mathrm{KL}\left(p\,||\,q\right)=\log\left\{\frac{\Gamma\left(a_{q}\right)\Gamma\left(b_{q}\right)}{\Gamma\left(a_{q}+b_{q}\right)}\right\}-\log\left\{\frac{\Gamma\left(a_{p}\right)\Gamma\left(b_{p}\right)}{\Gamma\left(a_{p}+b_{p}\right)}\right\}\\ +\left(a_{p}-a_{q}\right)\left\{\psi\left(a_{p}\right)-\psi\left(a_{p}+b_{p}\right)\right\}+\left(b_{p}-b_{q}\right)\left\{\psi\left(b_{p}\right)-\psi\left(a_{p}+b_{p}\right)\right\},$
@@
Since $\psi$ is strictly increasing, we observe that the right-hand side diverges as $a_{p}\to\infty$. Thus, the KL divergence between beta distributions is unbounded. The $h$-lifted KL divergence in contrast does not suffer from this problem, and does not require the restriction to $\mathcal{P}_{\alpha}$. This allows us to consider cases where $p,q\in\mathcal{P}$ are not bounded away from 00.
@@
$\mathrm{KL}_{h}\left(p\,||\,q\right)$ is bounded for all continuous densities $p,q\in\mathcal{P}$.
Proof.  Let $\tilde{p}=p+h$ and $\tilde{q}=q+h$. Since $h$ is positive, there exists some $\tilde{q}_{*}$ such that $\tilde{q}_{*}=\inf_{x\in\mathcal{X}}\left\{q(x)+h(x)\right\}>0$. Similarly, since $\mathcal{X}$ is compact, there exists some positive $\tilde{p}^{*}$ such that $0<\tilde{p}^{*}=\sup_{x\in\mathcal{X}}\left\{p(x)+h(x)\right\}<\infty$. Define $M=\sup_{x\in\mathcal{X}}\log\{\tilde{p}(x)/\tilde{q}(x)\}$. Then $M<\infty$, and
$\mathrm{KL}_{h}\left(p\,||\,q\right)=\int_{\mathcal{X}}\tilde{p}\log\frac{\tilde{p}}{\tilde{q}}\mathrm{d}\mu\leq\sup_{x\in\mathcal{X}}\log\frac{\tilde{p}}{\tilde{q}}\int_{\mathcal{X}}\tilde{p}\mathrm{d}\mu=2M<\infty.$
Let $L_{p}(f,g)$ denote the standard $L_{p}$ norm,
$L_{p}(f,g)=\left\{\int_{\mathcal{X}}\left\lvert f(x)-g(x)\right\rvert^{p}\mathrm{d}\mu(x)\right\}^{1/p}.$
As remarked previously, Klemelä (2007) established empirical risk bounds in terms of the $L_{2}$ norm distance. Following results from Meir and Zeevi (1997), characterizing the relationship between the KL divergence in terms of the $L_{2}$ norm distance, we establish the corresponding relationship between the $h$-lifted KL divergence and the $L_{2}$ norm distance.

### v2

@@ Abstract
We consider the problem of estimating probability density functions based on sample data, using a finite mixture of densities from some component class. To this end, we introduce the $h$-lifted Kullback–Leibler (KL) divergence as a generalization of the standard KL divergence and a criterion for conducting risk minimization. Under a compact support assumption, we prove an $\mathcal{O}(1/{\sqrt{n}})$ bound on the expected estimation error when using the $h$-lifted KL divergence, which extends the results of Rakhlin et al. (2005, ESAIM: Probability and Statistics, Vol. 9) and Li & Barron (1999, Advances in Neural Information Processing Systems, Vol. 12) to permit the risk bounding of density functions that are not strictly positive. We develop a procedure for the computation of the corresponding maximum $h$-lifted likelihood estimators ($h$-MLLEs) using the Majorization-Maximization framework and provide experimental results in support of our theoretical bounds.
@@ 1 Introduction
Let $\left(\Omega,\mathfrak{A},{\mathrm{\mathbf{P}}}\right)$ be an abstract probability space and let $X:\Omega\to\mathcal{X}$ be a random variable taking values in the measurable space $\left(\mathcal{X},\mathfrak{F}\right)$, where $\mathcal{X}$ is a compact metric space equipped with its Borel $\sigma$-algebra $\mathfrak{F}$. Suppose that we observe an independent and identically distributed (i.i.d.) sample of random variables $\mathbf{X}_{n}=\left(X_{i}\right)_{i\in\left[n\right]}$, where $\left[n\right]=\left\{1,\ldots,n\right\}$, and that each $X_{i}$ arises from the same data generating process as $X$, characterized by the probability measure $F\ll\mu$ on $\left(\mathcal{X},\mathfrak{F}\right)$, with density function $f=\mathrm{d}F/\mathrm{d}\mu$, for some $\sigma$-finite $\mu$. In this work, we are concerned with the estimating $f$ via a data dependent double-index sequence of estimators $\left(f_{k,n}\right)_{k,n\in\mathbb{N}}$, where
$f_{k,n}\in\mathcal{C}_{k}=\mathrm{co}_{k}\left(\mathcal{P}\right)=\left\{f_{k}\left(\cdot;\psi_{k}\right)=\sum_{j=1}^{k}\pi_{j}\varphi\left(\cdot;\theta_{j}\right)\mid\varphi\left(\cdot;\theta_{j}\right)\in\mathcal{P},\,\pi_{j}\geq 0,\,j\in\left[k\right],\sum_{j=1}^{k}\pi_{j}=1\right\},$
@@
$\mathcal{P}=\left\{\varphi\left(\cdot;\theta\right):\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}\mid\theta\in\Theta\subset\mathbb{R}^{d}\right\},$ (1)
$\psi_{k}=\left(\pi_{1},\dots,\pi_{k},\theta_{1},\dots,\theta_{k}\right)$, and $d\in\mathbb{N}$. To ensure the measurability and existence of various optima, we shall assume that $\varphi$ is Carathéodory in the sense that $\varphi\left(\cdot;\theta\right)$ is $\left(\mathcal{X},\mathfrak{F}\right)$-measurable for each $\theta\in\Theta$, and $\varphi\left(X;\cdot\right)$ is continuous for each $X\in\mathcal{X}$.
In the definition above, we can identify the set $\mathcal{C}_{k}=\mathrm{co}_{k}\left(\mathcal{P}\right)$ as the set of density functions that can be written as a convex combination of $k$ elements of $\mathcal{P}$, where $\mathcal{P}$ is often called the space of component density functions. We then interpret $\mathcal{C}_{k}$ as the class of $k$-component finite mixtures of densities of class $\mathcal{P}$, as studied, for example, by McLachlan & Peel (2004); Nguyen et al. (2020; 2022b).
@@ 1.1 Risk bounds for mixture density estimation
where $\left(p,q\right)\mapsto\ell\left(p,q\right)\in\mathbb{R}_{\geq 0}$ is a loss function on pairs of density functions. We define the density-to-class loss
@@
where $\mathrm{cl}(\cdot)$ is the closure. Here, we identify $\left(k,n\right)\mapsto\rho\left(k,n\right)$ as a characterization of the rate at which the left-hand side of (2) converges to zero as $k$ and $n$ increase. Our present work follows the research of Li & Barron (1999), Rakhlin et al. (2005) and Klemelä (2007) (see also Klemelä 2009, Ch. 19). In Li & Barron (1999) and Rakhlin et al. (2005), the authors consider the case where $\ell\left(p,q\right)$ is taken to be the Kullback–Leibler (KL) divergence
$\mathrm{KL}\left(p\,||\,q\right)=\int p\log\frac{p}{q}\mathrm{d}\mu$
@@
Under the assumption that $f,f_{k}\geq a$, for some $a>0$ and each $k\in\left[n\right]$ (i.e., strict positivity), Li & Barron (1999) obtained the bound
@@
for constants $c_{1},c_{2},c_{3}>0$, which was then improved by Rakhlin et al. (2005) who obtained the bound
@@
where $\lVert p\rVert_{2,\mu}^{2}=\int_{\mathcal{X}}\lvert p\rvert^{2}\mathrm{d}\mu$, for each $p\in L_{2}\left(\mu\right)$, and choose $f_{k,n}$ as minimizers of the $L_{2}\left(\mu\right)$ empirical risk, i.e., $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$, where
@@
$c_{1},c_{2}>0$, without the lower bound assumption on $f,f_{k}$ above, even permitting $\mathcal{X}$ to be unbounded. Via the main results of Naito & Eguchi (2013), the bound above can be generalized to the $U$-divergences, which includes the special $L_{2}(\mu)$ norm distance as a special case.
On the one hand, the sequence of MLEs required for the results of Li & Barron (1999) and Rakhlin et al. (2005) are typically computable, for example, via the usual expectation–maximization approach (cf. McLachlan & Peel 2004, Ch. 2). This contrasts with the computation of least-squares density estimators of form (3), which requires evaluations of the typically intractable integral expressions $\left\|f_{k}\left(\cdot;\psi_{k}\right)\right\|_{2}^{2}$. However, the least-squares approach of Klemelä (2007) permits the analysis using families $\mathcal{P}$ of usual interest, such as normal distributions and beta distributions, the latter of which being compactly supported but having densities that cannot be bounded away from zero without restrictions, and thus do not satisfy the regularity conditions of Li & Barron (1999) and Rakhlin et al. (2005).
@@ 1.2 Main contributions
We propose the following $h$-lifted KL divergence, as a generalization of the standard KL divergence to address the computationally tractable estimation of density functions which do not satisfy the regularity conditions of Li & Barron (1999) and Rakhlin et al. (2005). The use of the $h$-lifted KL divergence has the possibility to advance theories based on the standard KL divergence in statistical machine learning. To this end, let $h:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ be a function in $L_{1}(\mu)$, and define the $h$-lifted KL divergence by:
@@
Assume that $h$ is a probability density function, and let $\mathbf{Y}_{n}=\left(Y_{i}\right)_{i\in\left[n\right]}$ be a an i.i.d. sample, independent of $\mathbf{X}_{n}$, where each $Y_{i}:\Omega\rightarrow\mathcal{X}$ is a random variable with probability measure on $(\mathcal{X},\mathfrak{F})$, characterized by the density $h$ with respect to $\mu$. Then, for each $k$ and $n$, let $f_{k,n}$ be defined via the maximum $h$-lifted likelihood estimator ($h$-MLLE; see Appendix B for further discussion) $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$, where
@@
This result is a compromise between the works of Li & Barron (1999) and Rakhlin et al. (2005), and Klemelä (2007), as it applies to a broader space of component densities $\mathcal{P}$, and because the required $h$-MLLEs (5) can be efficiently computed via minorization–maximization (MM) algorithms (see e.g., Lange 2016). We shall discuss this assertion in Section 4.
@@ 1.3 Relevant literature
Our work largely follows the approach of Li & Barron (1999), which was extended upon by Rakhlin et al. (2005) and Klemelä (2007). All three texts use approaches based on the availability of greedy algorithms for maximizing convex functions with convex functional domains. In this work, we shall make use of the proof techniques of Zhang (2003). Related results in this direction can be found in DeVore & Temlyakov (2016) and Temlyakov (2016). Making the same boundedness assumption as Rakhlin et al. (2005), Dalalyan & Sebbar (2018) obtain refined oracle inequalities under the additional assumption that the class $\mathcal{P}$ is finite. Numerical implementations of greedy algorithms for estimating finite mixtures of Gaussian densities were studied by Vlassis & Likas (2002) and Verbeek et al. (2003).
The $h$-MLLE as an optimization objective can be compared to other similar modified likelihood estimators, such as the $L_{q}$ likelihood of Ferrari & Yang (2010) and Qin & Priebe (2013), the $\beta$-likelihood of Basu et al. (1998) and Fujisawa & Eguchi (2006), penalized likelihood estimators, such as maximum a posteriori estimators of Bayesian models, or $f$-separable Bregman distortion measures of Kobayashi & Watanabe (2024; 2021).
The practical computation of the $h$-MLLEs, (5), is made possible via the MM algorithm framework of Lange (2016), see also Hunter & Lange (2004), Wu & Lange (2010), and Nguyen (2017) for further details. Such algorithms have well-studied global convergence properties and can be modified for mini-batch and stochastic settings (see, e.g., Razaviyayn et al., 2013 and Nguyen et al., 2022a).
A related and popular setting of investigations is that of model selection, where the objects of interest are single-index sequences $\left(f_{k_{n},n}\right)_{n\in\mathbb{N}}$, and where the aim is to obtain finite-sample bounds for losses of the form $\ell\left(f_{k_{n},n},f\right)$, where each $k_{n}\in\mathbb{N}$ is a data dependent function, often obtained by optimizing some penalized loss criterion, as described in Massart (2007), Koltchinskii (2011, Ch. 6), and Giraud (2021, Ch. 2). In the context of finite mixtures, examples of such analyses can be found in the works of Maugis & Michel (2011) and Maugis-Rabusseau & Michel (2013). A comprehensive bibliography of model selection results for finite mixtures and related statistical models can be found in Nguyen et al. (2022c).
@@ 1.4 Organization of paper
The manuscript is organized as follows. In Section 2, we formally define the $h$-lifted KL divergence as a Bregman divergence and establish several of its properties. In Section 3, we present new risk bounds for the $h$-lifted KL divergence of the form (2). In Section 4, we discuss the computation of the $h$-lifted likelihood estimator in the form of (5), followed by empirical results illustrating the convergence of (2) with respect to both $k$ and $n$. Additional insights and technical results are provided in the Appendices at the end of the manuscript.
@@ 2.1 $\textrm{KL}_{h}$ as a Bregman divergence
The properties for Bregman divergences between scalars can be extended to density functions and other functional spaces, as established in Frigyik et al. (2008) and Stummer & Vajda (2012), for example. We also direct the interested reader to the works of Pardo (2006), Basu et al. (2011), and Amari (2016).
@@
The $h$-lifted KL divergence $\mathrm{KL}_{h}$ as a Bregman divergence is generated by the function $\phi(u)=(u+h)\log(u+h)-(u+h)+1$. This assertion is demonstrated in Appendix C.1.
@@ 2.2 Advantages of the $h$-lifted KL divergence
When the standard KL divergence is employed in the density estimation problem, it is common to restrict consideration of density functions to those bounded away from zero by some positive constant. That is, one typically considers the smaller class of so-called admissible target densities $\mathcal{P}_{\alpha}\subset\mathcal{P}$ (cf. Meir & Zeevi, 1997), where
@@
$\mathrm{KL}\left(p\,||\,q\right)$ $=\log\left\{\frac{\Gamma\left(a_{q}\right)\Gamma\left(b_{q}\right)}{\Gamma\left(a_{q}+b_{q}\right)}\right\}-\log\left\{\frac{\Gamma\left(a_{p}\right)\Gamma\left(b_{p}\right)}{\Gamma\left(a_{p}+b_{p}\right)}\right\}$
$\quad+\left(a_{p}-a_{q}\right)\left\{\psi\left(a_{p}\right)-\psi\left(a_{p}+b_{p}\right)\right\}+\left(b_{p}-b_{q}\right)\left\{\psi\left(b_{p}\right)-\psi\left(a_{p}+b_{p}\right)\right\},$
@@
Since $\psi$ is strictly increasing, we observe that the right-hand side diverges as $a_{p}\to\infty$. Thus, the KL divergence between beta distributions is unbounded. The $h$-lifted KL divergence in contrast does not suffer from this problem, and does not require the restriction to $\mathcal{P}_{\alpha}$. This allows us to consider cases where $p,q\in\mathcal{P}$ are not bounded away from 00, as per the following result.
@@
Let $\mathcal{P}$ be defined as in (1). $\mathrm{KL}_{h}\left(f\,||\,g\right)$ is bounded for all continuous densities $f,g\in\mathcal{P}$.
Proof.
See Appendix C.2. ∎
Let $L_{p}(f,g)$ denote the standard $L_{p}$-norm, $L_{p}(f,g)=\left\{\int_{\mathcal{X}}\left\lvert f(x)-g(x)\right\rvert^{p}\mathrm{d}\mu(x)\right\}^{1/p}$. As remarked previously, Klemelä (2007) established empirical risk bounds in terms of the $L_{2}$-norm distance. Following results from Meir & Zeevi (1997) characterizing the relationship between the KL divergence in terms of the $L_{2}$-norm distance, in Proposition 3 we establish the corresponding relationship between the $h$-lifted KL divergence and the $L_{2}$-norm distance, along with a relationship between the $h$-lifted KL divergence and the $L_{1}$-norm distance.